# Generative Video Compression: Towards 0.01% Compression Rate for Video Transmission

### v1

@@ 3 Results
HEVC Sullivan et al. (2012) 0.271
Ours 0.214
To further validate its practical utility, we apply the reconstruction results of the model to the downstream task: video object segmentation (VOS) on DAVIS2017 (Pont-Tuset et al., 2018). We evaluate performance using the Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, their average ($\mathcal{J}$$\mathcal{F}$), and contour recall ($\mathcal{F}$-Recall). As shown in Table 2, our method achieves highly competitive performance. This indicates that even at low bitrates, our approach can preserve correct semantic transmission.
@@
Although the miniaturized model incurs some loss in visual quality and bandwidth efficiency compared to its full-scale counterpart, it still maintains competitively high perceptual quality, as illustrated in Figure 3, where the shown video sequence achieves LPIPS of 0.342. These results collectively demonstrate that our miniaturized model achieves an effective balance between computational efficiency and visual quality in practical deployment scenarios.

### v2

@@ 3 Results
HEVC Sullivan et al. (2012) 0.278
Ours 0.180
To further validate its practical utility, we apply the reconstruction results of the model to the downstream task: video object segmentation (VOS) on DAVIS2017 (Pont-Tuset et al., 2017). We evaluate performance using the Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, their average ($\mathcal{J}$$\mathcal{F}$), and contour recall ($\mathcal{F}$-Recall). As shown in Table 2, our method achieves highly competitive performance. This indicates that even at low bitrates, our approach can preserve correct semantic transmission.
@@
Although the miniaturized model incurs some loss in visual quality and bandwidth efficiency compared to its full-scale counterpart, it still maintains competitively high perceptual quality, as illustrated in Figure 3, where the shown video sequence achieves LPIPS of 0.273. These results collectively demonstrate that our miniaturized model achieves an effective balance between computational efficiency and visual quality in practical deployment scenarios.